# Equal contribution

### v3

@@ 4.4 Long Group Diffusion Sampling
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation, $\boldsymbol{{z}}_{T}^{[151:225]}$ is a random noise, and $\widehat{\boldsymbol{{z}}}_{T}^{[75:225]}$ is the final input noise. Compared to the existing method, this approach, which constrains the input noise, introduces less randomness and provides cleaner positional constraints during generation, thus effectively improving the consistency of the transitions in long-term generation. We provide further analysis in the ablation study section, demonstrating the effectiveness of our approach.
@@ 5.1 Experimental Settings
Implementation Details. Our model employs a diffusion-based architecture comprising the Group Dance Decoder (GDD) and the Footwork Adaptor (FA). We set the hidden size of all components in our framework to $d=512$. In GDD, a linear layer first transforms the 151-dimensional input motion into a 512-dimensional hidden representation. This representation is then processed by the Fusion Projection, which consists of a three-layer MLP with ReLU activation. Next, the Sequence Decoder, composed of $M=8$ stacked layers and equipped with eight attention heads, extracts sequential information from the data. Finally, an output linear layer maps the hidden features back to 151-dimensional raw motion data. For footwork modulation, we construct the FA using $W=3$ stacked Concat Squash Linear layers. We set the weighting coefficients as follows: $\lambda_{\text{D}}=\lambda_{\text{Sim}}=0.636$, $\lambda_{\text{vel}}=2.964$, $\lambda_{\text{FK}}=0.646$, and $\lambda_{\text{con}}=10.942$. We used batch sizes of 72, 53, 40, and 23 per GPU for 2, 3, 4, and 5 dancers, respectively. We using Adam for optimization with a learning rate of $5\times 10^{-5}$.
@@ 5.2 Qualitative Visual Comparison
Table 1: Quantitative comparison with the baselines.
@@
00footnotetext: $\uparrow$ means higher is better, $\downarrow$ means lower is better. The best results are highlighted in bold, the second best results are underlined.
@@ 5.3 Quantitative Comparison
Table 2: Quantitative comparison with the baselines for generating 2-5 dancers.
@@
00footnotetext: $\uparrow$ means higher is better, $\downarrow$ means lower is better.
Figure 8: User study based on four criteria: motion realism, music-motion correlation, formation aesthetics, and harmony of dancers. Our model has garnered greater user favor, showcasing our superiority in aesthetic appeal.
@@
User study Our user study results are shown in Figure 8. The results show that the visual effects generated by our model are the most favored by users. This is because TCDiff++ effectively mitigates significant visual flaws such as collisions and foot sliding, while producing more consistent footwork movements and resulting in more harmonious dancer formations.
Table 3: Baseline comparison in a long-duration setting.
@@ 5.4 Ablation Study
To evaluate the effectiveness of our designed module under the long-duration scenario (720 frames), the results are shown in Table 4. We ranked the performance of different ablated models from worst to best to provide an intuitive ordering of performance improvement. The results show that the model achieves the best performance when all modules are applied simultaneously (Full). Overall, all modules have contributed to improving the model’s performance on group metrics, thereby demonstrating their effectiveness in enhancing the harmony of group dance. This improvement is attributed to their varying degrees of mitigation of multi-dancer collisions and foot sliding, leading to more realistic dance movements. Notably, our proposed Fusion Projection (FP) and Long Group Diffusion Sampling (LGDS) significantly reducing abrupt swapping (PFC and TIF), as shown in Figure 9 and 10 enhancing the visual quality of the generated results. Among them, the simple-yet-effective FP module contributes the most to improving the performance of the end-to-end model, effectively mitigating the adverse effects of dancer ambiguity. Meanwhile, the LGDS strategy enhances the model’s performance in extending group dance motion sequences within the diffusion framework in a training-free manner, effectively reducing the occurrence of abrupt swapping. In mitigating foot sliding, the Footwork Adaptor and Swap Mode modules play a significant role by enhancing the model’s ability to capture sequential dependencies and incorporating additional spatial information. In alleviating multi-dancer collision, the Dancer Positioning Loss (DPE) and distance-consistency loss $\mathcal{L}_{D}$ are particularly effective, as they amplify feature distinctions among dancers and introduce spatial constraints.
For simplicity, we use the following abbreviations throughout the subsequent sections: Fusion Projection (FP), Footwork Adaptor (FA), Sequence Decoder (SD), Dancer Positioning Embedding (DPE), Swap Mode (SM), Long Group Dance Sampling (LGDS) and $\mathcal{L}_{D}$ for distance-consistency loss.
Table 4: Ablation study results.
@@
w/o FP 30.75 79.14 0.23 43.69 17.48 0.21 1.89
w/o FA 25.37 80.75 0.15 33.27 12.75 0.22 2.34
w/o SD 25.75 80.72 0.17 28.75 14.24 0.25 1.52
w/o DPE 20.97 80.91 0.18 40.75 10.52 0.23 1.58
w/o SM 16.71 81.60 0.17 25.97 15.75 0.23 1.75
w/o LGDS 15.72 80.34 0.18 36.23 10.52 0.22 1.92
w/o $\mathcal{L}_{D}$ 18.72 80.91 0.18 24.34 15.75 0.23 1.87
Full 14.67 81.64 0.15 20.37 16.19 0.23 1.53
Effectiveness of FP. With FP, the model demonstrates significant improvements across both multi-dancer and single-dancer metrics. GMR decreases from 30.75 to 14.67, indicating enhanced group formation quality, while GMC increases from 79.14 to 81.64, reflecting stronger correlation among dancers. Additionally, TIF is reduced from 0.23 to 0.15, suggesting fewer dancer collisions and improved formation organization. With FP, the model reduces random abrupt swapping and unrealistic formations, resulting in more structured and coherent dance generation. This refinement leads to a slight decrease in Div from 17.48 to 16.19, while significantly lowering FID from 43.69 to 20.37, thereby enhancing the realism of the generated dances. Lastly, PFC decreases from 1.89 to 1.53, improving the physical plausibility of footwork. These results confirm that FP plays a crucial role in refining both structural coherence and visual quality in dance generation. Figure 10 illustrates that integrating FP effectively reduces dancer overlap and abrupt swapping. The coordinate trajectories reveal a more reasonable dancer formation, further improving the visual quality of the generated results. This is because FP amplifies the differences between dancers in high-dimensional space, which effectively helps reduce dancer ambiguity.
@@
Effectiveness of FA. Our FA markedly enhances the model’s PFC value, and even enhances model’s performance in group dance metrics and FID. Notably, adding FA reduces PFC from 2.34 to 1.53, enhancing the physical plausibility of footwork. With FA, GMR decreases from 25.37 to 14.67, indicating improved group formation quality, while GMC increases from 80.75 to 81.64, enhancing dancer correlation. For single-dancer evaluation, FA significantly reduces FID from 33.27 to 20.37, enhancing the similarity to real dance sequences by ensuring more realistic footwork. Moreover, Div increases from 12.75 to 16.19, indicating greater motion variety. This is because FA captures the connection between positional shifts and footwork, leading to more realistic dance footwork.
Effectiveness of SD. For group-dance evaluation, GMR decreases from 25.75 to 14.67, indicating a significant improvement in group formation realism. GMC increases slightly from 80.72 to 81.64, suggesting better dancer coordination. TIF also decreases from 0.17 to 0.15, reflecting fewer dancer collisions. For single-dance evaluation, FID improves from 28.75 to 20.37, indicating better alignment with real dance sequences. Div increases from 14.24 to 16.19, showing more motion variety in the full model. Our SD leverages the SSM architecture to enhance selection in sequential data, therefore it becomes easier to generate movements with rich expressiveness and less susceptible to noise in long-duration scenario. Although the increased variety (Div) introduces some redundant movements that are slightly misaligned with the musical rhythm (lower MMC), it significantly enhances the model’s visual performance, as reflected in the improved group-dance metrics.
Effectiveness of DPE. When compared to the full model, the results show several improvements. GMR increases from 20.97 to 14.67, indicating a significant enhancement in group formation realism with the inclusion of DPE. GMC remains stable, with a slight increase from 80.91 to 81.64, reflecting improved coordination among dancers. TIF decreases from 0.18 to 0.15, showing a reduction in dancer collisions. In terms of single-dance evaluation, FID improves from 40.75 to 20.37, reflecting a better alignment of generated dances with real sequences. Div increases from 10.52 to 16.19, indicating greater motion variety in the full model. MMC remains unchanged at 0.23, indicating consistent synchronization with the music’s rhythm. PFC slightly decreases from 1.58 to 1.53, showing a slight improvement in footwork plausibility. DPE effectively mitigates multi-dancer collisions (lower TIF), as it incorporates the left-right spatial information of dancers, which helps maintain their relative positioning and enhances the overall performance of the model (improved group-dance metrics and FID).
Effectiveness of SM. When compared to the full model, with SM, the GMR increases from 16.71 to 14.67, showing an improvement in group formation realism. TIF decreases slightly from 0.17 to 0.15, suggesting fewer dancer collisions. For single-dance evaluation, FID improves from 25.97 to 20.37, indicating better alignment with real dance sequences. Div decreases slightly from 15.75 to 16.19, reflecting a marginal increase in motion variety. PFC decreases slightly from 1.75 to 1.53, reflecting a slight improvement in footwork plausibility. The improvements can be attributed to SM’s role in reducing spatial uncertainty caused by swap actions. By providing the dancers’ start and end positions, SM helps mitigate the uncertainty during spatial changes (e.g., swaps), resulting in better group coherence, more realistic footwork, and enhanced overall performance.
Effectiveness of LGDS. LGDS primarily mitigates abrupt swapping during long-duration extension. Although this transient phenomenon has little impact on quantitative metrics due to its short duration, it significantly degrades visual quality. Thus, while LGDS shows less improvement than FP in metrics, it plays a crucial role in enhancing visual perception. LGDS significantly reduces the occurrence of abrupt swapping in long-term generation (lower TIF) and can be applied training-free during the inference stage, leading to a substantial improvement in model performance. This is because LGDS enforces consistency during the input phase, thereby reducing randomness and ensuring cleaner positional information during generation, which in turn minimizes abrupt swaps. Figure 9 demonstrates that the incorporation of LGDS effectively mitigates abrupt dancer swapping in long-term generation, ensuring greater spatial consistency during sequence extension and significantly enhancing the visual coherence of group dance.
Effectiveness of $\mathcal{L}_{D}$. By emphasizing proper spatial relationships, $\mathcal{L}_{D}$ helps maintain optimal distances between dancers. This results in more natural and coherent group formations. Consequently, the model achieves better overall performance in group-dance metrics.

### v4

@@ 4.4 Long Group Diffusion Sampling
where $\cdot$$\cdot$] denotes the concatenation operation, $\boldsymbol{{z}}_{T}^{[151:225]}$ is a random noise, and $\widehat{\boldsymbol{{z}}}_{T}^{[75:225]}$ is the final input noise. Compared to the existing method, this approach, which constrains the input noise, introduces less randomness and provides cleaner positional constraints during generation, thus effectively improving the consistency of the transitions in long-term generation. We provide further analysis in the ablation study section, demonstrating the effectiveness of our approach.
@@ 5.1 Experimental Settings
Implementation Details. Our model employs a diffusion-based architecture comprising the Group Dance Decoder (GDD) and the Footwork Adaptor (FA). We set the hidden size of all components in our framework to $d=512$. In GDD, a linear layer first transforms the 151-dimensional input motion into a 512-dimensional hidden representation. This representation is then processed by the Fusion Projection, which consists of a three-layer MLP with ReLU activation. Next, the Sequence Decoder, composed of $M=8$ stacked layers and equipped with eight attention heads, extracts sequential information from the data. Finally, an output linear layer maps the hidden features back to 151-dimensional raw motion data. For footwork modulation, we construct the FA using $W=3$ stacked Concat Squash Linear layers. We set the weighting coefficients as follows: $\lambda_{\text{Sim}}=0.636$, $\lambda_{\text{vel}}=2.964$, $\lambda_{\text{FK}}=0.646,\lambda_{\text{D}}=100$, and $\lambda_{\text{con}}=10.942$. We used batch sizes of 72, 53, 35, and 23 per GPU for 2, 3, 4, and 5 dancers, respectively. We use Adam for optimization with a learning rate of $5\times 10^{-5}$.
@@ 5.2 Qualitative Visual Comparison
Table 1: Quantitative comparison with the baselines.00footnotetext: $\uparrow$ means higher is better, $\downarrow$ means lower is better. The best results are highlighted in bold, the second best results are underlined.
@@ 5.3 Quantitative Comparison
Table 2: Quantitative comparison with the baselines for generating 2-5 dancers.00footnotetext: $\uparrow$ means higher is better, $\downarrow$ means lower is better.
@@
Figure 8: User Study Results. Our model received higher user ratings, demonstrating its superior aesthetic appeal compared to other methods.
@@
User study Our user study results are presented in Figure 8. The study involved participants primarily from a younger demographic (ages 18–60, mostly in their 20s and 30s), aligning with the target user group of our application scenario. Regarding the experimental procedure, we anonymized the outputs generated by different methods and randomly mixed them before presenting them to participants. They were then asked to rate the outputs on a scale from 1 to 10 based on four criteria: motion realism, music-motion correlation, formation aesthetics, and dancer harmony. The final voting results were normalized to a $[0,1]$ range for consistent comparison. To ensure fairness, none of the samples from the previous conference version were reused, ensuring that all outputs remained anonymous to users. As a result, slight differences in scores may occur. Nevertheless, the results indicate that the visual effects generated by our model are the most favored by users, as TCDiff++ effectively reduces visual artifacts such as collisions and foot sliding, while producing more consistent footwork and harmonious group formations.
Table 3: Baseline comparison in a long-duration setting. 00footnotetext: $\uparrow$ means higher is better, $\downarrow$ means lower is better. The best results are highlighted in bold.
@@ 5.4 Ablation Study
Table 4: Ablation study evaluating the incremental contributions of each component in TCDiff++. The best results are highlighted in bold, the second best results are underlined
@@
Base 67.24 57.65 0.38 55.40 19.97 0.23 3.87
+FP 50.12 77.80 0.31 44.31 16.90 0.20 3.01
+FP+DPE 44.38 78.22 0.28 40.75 16.82 0.22 3.08
[…]
@@
We conduct a step-by-step ablation to evaluate the incremental contributions of each module. We provide visual comparisons in our demo video, available on our project page, for a more intuitive understanding of the improvements. All modules bring clear and cumulative improvements, both quantitatively and visually. Our ablation study not only presents clear quantitative improvements but also highlights how each module addresses key visual challenges in long-duration group dance generation. Detailed results are shown in Table 4. Note that the high diversity score of the Base model is largely due to irregular abrupt swaps, which do not indicate good visual quality. These severe artifacts are significantly reduced after introducing $\mathcal{L}_{D}$. Below is a concise analysis of each component:
Baseline Model: The baseline model [50] exhibits significant motion instability, with frequent abrupt position swaps and severe jitter in long sequences. This results in high GMR (67.24), poor group coherence, with a GMC score of 57.65, and high TIF (0.38), indicating that dancers often collide or lose spatial consistency. Note that the high diversity score is largely due to irregular abrupt swaps, which do not indicate good visual quality.
Impact of Fusion Projection (FP): FP enhances the model’s ability to distinguish between dancers by projecting their features into a shared high-dimensional space. This leads to a clear improvement in group motion quality, reflected in a reduction of GMR from 67.24 to 50.12 and an increase in GMC from 57.65 to 77.80. Additionally, it helps slightly reduce foot collisions, as indicated by the PFC decrease from 3.87 to 3.01.
Impact of Dancer Positioning Embedding (DPE): The introduction of DPE stabilizes the dancers’ trajectories by embedding both temporal and identity information. This reduces jitter, improving the overall visual stability, as shown by further reductions in GMR (50.12 $\rightarrow$ 44.38) and TIF (0.31 $\rightarrow$ 0.28). However, PFC slightly increases from 3.01 to 3.08, suggesting that while overall group structure improves, foot sliding issues persist without additional motion constraints.
Impact of Distance-Consistency Loss ($\mathcal{L}_{D}$): $\mathcal{L}_{D}$ imposes global spatial constraints that maintain consistent inter-dancer spacing. This significantly reduces collisions, driving GMR down from 44.38 to 25.76 and TIF from 0.28 to 0.19. It also lowers PFC (3.08 $\rightarrow$ 2.54), indicating less foot sliding. Nevertheless, due to the probabilistic nature of diffusion, occasional abrupt swaps can still occur in long sequences.
Impact of Swap Mode Embedding (SM): SM encodes dancer positional information at segment boundaries, effectively reducing confusion in spatial transitions. This results in improved formation and decreased TIF (0.19 $\rightarrow$ 0.17). GMC also improves slightly (81.01 $\rightarrow$ 81.45), confirming better temporal coordination during swaps.
Impact of Long Group Diffusion Sampling (LGDS): LGDS enhances long-term temporal consistency by conditioning each segment on previously generated frames. This stabilizes continuity, particularly in long sequences. The main visual improvement lies in the further reduction of severe abrupt swaps. GMR slightly improves (23.93 $\rightarrow$ 23.52), and TIF decreases to 0.16, indicating fewer inter-dancer collisions. PFC also drops from 2.60 to 2.48, reinforcing better lower-body coherence.
Impact of Footwork Adaptor (FA): FA focuses on refining lower-body movement by aligning foot trajectories with position shifts. This leads to a significant drop in PFC (2.48 $\rightarrow$ 1.49), confirming reduced foot sliding and improving the physical realism of foot placement. MMC improves from 0.22 to 0.24, indicating better alignment with music rhythms, while other metrics (e.g., FID drops from 27.50 to 24.75) also suggest enhanced visual quality.
Impact of Sequence Decoder (SD): The SD replaces the Transformer with a State Space Model, enabling more efficient and coherent long-sequence modeling. This results in the largest gains: GMR drops significantly from 23.95 to 14.67, FID improves from 24.75 to 20.37, and Div increases from 14.88 to 16.19. Despite a slight increase in PFC (1.49 $\rightarrow$ 1.53), the generated dances exhibit more fluid, expressive, and temporally consistent motion.
In conclusion, each component in our model addresses specific visual issues, from jitter and foot sliding to positional swaps and long-duration coherence. FP and DPE improve spatial awareness; $\mathcal{L}_{D}$, SM, and LGDS enhance group formation and reduce swapping; FA reduces foot sliding and enhances physical realism; and SD boosts long-term coherence and expressiveness. These improvements collectively yield the best performance across all group- and single-dancer metrics, validating the design of our modular architecture.
6 Limitation
Due to the early stage of group dance research, our work still has limitations. First, our model lacks support for multimodal or user-controllable inputs. It focuses solely on the basic cross-modal generation task (e.g., generating dance from music), without incorporating any additional control signals. While this provides a necessary foundation, real-world applications often require richer forms of guidance, such as textual descriptions, motion keyframes, or genre-specific conditioning, to better reflect user intent. Given the ongoing challenges in visual quality and temporal consistency, we prioritize solving the core generation problem in this work. We leave multimodal and interactive control for future exploration, which we believe is essential for practical deployment, especially in complex group dance scenarios. Second, our model shows limited capability in learning effective swap patterns. This arises partly because modeling swap actions is still in its infancy and, more crucially, because existing datasets lack sufficient samples and explicit annotations of swap behaviors. As a result, performance on swap-related generation remains constrained. We believe future work should improve both model architectures and data resources to better capture and generate swap behaviors in group choreography.